Assume that a set of test scores is normally distributed with a mean of 100 and a standard deviation of 25. Use the​ 68-95-99.7 rule to find the following quantities.
Answer #1

Please don't hesitate to give a "thumbs up" in case you're satisfied with the answer

Mean = 100

Stdev = 25

68%-95%-99.7% values are 1, 2 and 3 deviations away from mean

Therefore,

1 deviation more than mean is 100+25 = 125

2 deviations more than mean is 100+2*25 = 150

3 deviations more than mean is 100+3*25 = 175

1 deviation less than mean is 100-25 = 75

2 deviations less than mean is 100-2*25 = 50

3 deviations less than mean is 100-3*25 = 25